# references_lipschitz_polytopes.bib

### v2

@@ 6.1. Tree-like metrics
The set of tree-like metrics is a subfan of $\mathscr{F}_{n}$. In particular, the set of phylogenetic trees on $n$ leaves is a union of (relatively open) cells of $\mathscr{F}_{n}$.
@@
Since each $\mathscr{C}_{\pi_{A}}$ is a subfan of $\mathscr{F}_{n}$, it follows that the set of all tree-like metrics is a subfan of $\mathscr{F}_{n}$ as well. ∎
@@ 6.2. Circular-decomposable metrics
Accordingly, the circular decomposable metrics form a full-dimensional subfan $\mathscr{K}_{n}$ of $\mathscr{F}_{n}$.
@@
Equation 6.3 implies immediately that $\overline{\mathscr{K}}_{n}$ is indeed a subfan of ¯$\overline{\mathscr{F}}_{n}$. The claim about simpliciality and dimension follows from Lemma 6.3. ∎
@@ 6.3. Split decomposition
Note that the notion of ”generic” metric discussed in [BandeltDress] is not the same as the one of Definition 1.1. Both metrics in Example 6.13 are generic in the sense of our Definition 1.1, but only $\rho_{2}$ is generic in the sense of [BandeltDress], while $\rho_{1}$ is ”degenerate” in the sense of Bandelt and Dress because the metrics used in the canonical decomposition of $\rho_{1}$ are a strict subset of those for $\rho_{2}$ (indeed $\rho_{2}$ has an additional nonzero contribution from a split-prime metric).
@@ 6.4. Consistent totally split-decomposable metrics
The set of metrics that satisfy the six-point condition Section 6.4 is a subfan of ¯$\overline{\mathscr{F}}_{n}$.
@@
This proves that the set of $\rho$ that satisfy the six-point condition is obtained uniting and intersecting sets of (closed) chambers, hence it is a union of faces of the arrangement ¯$\overline{\mathscr{F}}_{n}$. ∎
@@ 7. The $f$-vector in the strict case
If $r_{2}(\rho)=0$ then the third case never enters and so exactly one of the other alternatives is realized. Moreover, in this setting the admissible graphs obtained in the first two cases are face-defining according to Lemma 7.4, and we obtain a total of $\frac{(n+1)!}{4(n-3)!}$ admissible graphs of type (a). Thus, when $\rho$ is not contained in any member of $\mathcal{W}_{l}^{2}$ the proposition holds.
@@
The first case is trivial, the case of the ground set appearing once is described above and in the last case, we subtract all $6$ faces ($2$ per cycle) from the generic count. Doing this iteratively for each quadruple gives the result. e therefore subtract them from the count. Faces of dimension $1$ which correspond to paths on three vertices are not affected. ∎
@@ 8.1. The case $n=4$
Figure 7. Wasserstein arrangement for $n=4$. Each hyperplane corresponds to a cycle as described in Remark 3.6 and each chamber has one set of oriented edges of each cycle that is admissible. We obtain the admissible graphs of a neighboring chamber by applying the permutation specified on the hyperplane separating the chambers. The edge pairs in the gray area are the ones corresponding to the horizontal hyperplane.
@@ 8.2. The case $n=5$
For $6$-point metrics we employed two approaches. First, we computed the number of chambers of the arrangement $\mathcal{W}_{6}$ using the julia module CountingChambers.jl [BEK23]. This is an arrangement in $\mathbb{R}^{15}$ with a $6$-dimensional lineality space consisting of $105$ hyperplanes, see Proposition 3.7. We found that this arrangement has $6,677,863,200$ chambers (this agrees by the above discussion with the number of combinatorial types of generic KRW polytopes).
As explained in Section 4, the generic KRW polytopes correspond exactly to certain regular triangulations of $RP^{0}_{n}$, the (full) root polytope together with the origin, see Proposition 4.3. Jörg Rambau computed all symmetric regular triangulations of $RP^{0}_{6}$ up to symmetry using the latest version of his software TOPCOM [Ram23]. There are exactly $25,224$ such triangulations. This number therefore agrees with the number of combinatorial types of KRW polytopes of $6$-point metrics up to the action of the symmetry group $S_{6}$. We record both statements in the following theorem.
@@
There are exactly $6,677,863,200$ combinatorial types of generic KRW polytopes corresponding to $6$-point metrics. Under the action of the symmetric group $S_{6}$ these polytopes come in $25,224$ orbits.

### v3

@@ 6.1. Tree-like metrics
The set of tree-like metrics is a subfan of $\mathscr{F}_{n}$ and $\mathscr{T}_{n}$. In particular, the set of phylogenetic trees on $n$ leaves is a union of (relatively open) cells of $\mathscr{F}_{n}$.
@@
Since each $\mathscr{C}_{\pi_{A}}$ is a subfan of $\mathscr{F}_{n}$, it follows that the set of all tree-like metrics is a subfan of $\mathscr{F}_{n}$ as well. According to Proposition 3.13.(1) these hyperplanes are all part of the type fan $\mathscr{T}_{n}$. Hence, the set of tree-like metrics is also a subfan of $\mathscr{T}_{n}$. ∎
@@ 6.2. Circular-decomposable metrics
Accordingly, the circular decomposable metrics form a full-dimensional subfan $\mathscr{K}_{n}$ of $\mathscr{F}_{n}$. Moreover, $\mathscr{K}_{n}$ is also a full-dimensional subfan of the type fan $\mathscr{T}_{n}$.
@@
Equation 6.3 implies immediately that $\overline{\mathscr{K}}_{n}$ is indeed a subfan of ¯$\overline{\mathscr{F}}_{n}$. The claim about simpliciality and dimension follows from Lemma 6.3. As above, the statement on the type fan follows from Proposition 3.13. ∎
@@ 6.3. Split decomposition
Note that the notion of ”generic” metric discussed in [BandeltDress] is not the same as the one of Definition 1.1.
Both metrics in Example 6.13 are generic in the sense of our Definition 1.1, but only $\rho_{2}$ is generic in the sense of [BandeltDress], while $\rho_{1}$ is ”degenerate” in the sense of Bandelt and Dress because the metrics used in the canonical decomposition of $\rho_{1}$ are a strict subset of those for $\rho_{2}$ (indeed $\rho_{2}$ has an additional nonzero contribution from a split-prime metric).
@@ 6.4. Consistent totally split-decomposable metrics
The set of metrics that satisfy the six-point condition Section 6.4 is a subfan of ¯$\overline{\mathscr{F}}_{n}$ and of the type fan $\mathscr{T}_{n}$.
@@
This proves that the set of $\rho$ that satisfy the six-point condition is obtained uniting and intersecting sets of (closed) chambers, hence it is a union of faces of the arrangement ¯$\overline{\mathscr{F}}_{n}$. The statement on the type fan follows again from Proposition 3.13. ∎
@@ 7. The $f$-vector in the strict case
If $r_{2}(\rho)=0$ then the third case never enters and so exactly one of the other alternatives is realised. Moreover, in this setting the admissible graphs obtained in the first two cases are face-defining according to Lemma 7.4, and we obtain a total of $\frac{(n+1)!}{4(n-3)!}$ admissible graphs of type (a). Thus, when $\rho$ is not contained in any member of $\mathcal{W}_{l}^{2}$ the proposition holds.
@@
The first case is trivial, the case of the ground set appearing once is described above and in the last case, we subtract all $6$ faces ($2$ per cycle) from the generic count. Doing this iteratively for each quadruple gives the result. ∎
@@ 8.1. The case $n=4$
Figure 7. Wasserstein arrangement for $n=4$. Each hyperplane corresponds to a cycle as described in Remark 3.6 and each chamber has one set of oriented edges of each cycle that is admissible. We obtain the admissible graphs of a neighboring chamber by applying the permutation specified on the hyperplane separating the chambers. The edge pairs in the grey area are the ones corresponding to the horizontal hyperplane.
@@ 8.2. The case $n=5$
For $6$-point metrics we employed two approaches. First, we computed the number of chambers of the arrangement $\mathcal{W}_{6}$ using the julia module CountingChambers.jl [BEK23]. This is an arrangement in $\mathbb{R}^{15}$ consisting of $105$ hyperplanes with a $6$-dimensional lineality space, see Proposition 3.7. We found that $\mathcal{W}_{6}$ has $6,677,863,200$ chambers.
As explained in Section 4, the generic KRW polytopes correspond exactly to certain regular triangulations of $RP^{0}_{n}$, the (full) root polytope together with the origin, see Proposition 4.3. Jörg Rambau computed there there are $17,695,320$ symmetric central regular triangulations of $RP^{0}_{6}$ using the latest version of his software TOPCOM [Ram23]. This count agrees by the above discussion with the number of labeled combinatorial types of generic KRW polytopes. Up to symmetry, there are exactly $25,224$ such triangulations. This number therefore agrees with the number of KRW polytopes of $6$-point metrics up to the action of the symmetry group $S_{6}$. We record both statements in the following theorem.
@@
There are exactly $17,695,320$ labeled combinatorial types of generic KRW polytopes corresponding to $6$-point metrics. Under the action of the symmetric group $S_{6}$ these polytopes come in $25,224$ orbits.